Grade 11General Physics

v=under root d /p is incorrect and fined correct relation dimensionally where v is speed od sound and d is density and p is pressure

yashasvi srivastava
9 Years agoGrade 11

1 Answer

Anonymous
9 Years ago
[v]=[LT-1]  [d]=[ML-3]    [P]=[ML-1T-2]
 
Let [v]=[P]x[d]y
        
             [LT-1]=[ML-1T-2]x   [ML-3]y ​= [Mx+yL-x-3yT-2x]
 
x+y=0     -x-3y=1       -2x=-1
 
x=1/2   y=-1/2
 
 
 
V=(P/d)1/2 is the correct relation.After first finding the dimensions of each physical quantity we try to equate the  dimensions and use simultaneous equation.